How would you make a 350 ml solution of 2% Glucose (MW= 180.16) in water? A) 2 grams of glucose into 300 ml of water. Bring up to 350 ml total volume. B) Add 350 ml of distilled water to a beaker. Add 7 grams of glucose. C) 7 grams of glucose into 300 ml of water. Bring up to 350 ml total volume D) Add 350 ml of distilled water to a beaker. Add 20 grams glucose.
Answer #1

When a dry chemical is used it is mixed as dry mass per volume where g/100 ml is percent concentration.

Thus,

to make 350 ml solution of 2% glucose in water :

2g * ( 350 ml /100 ml) = 2* 3.5 g = 7 g

Thus, 7 grams of glucose should be added to the beaker and the volume should be made to 350 ml.

Thus option C is the most appropriate answer.